# Strategic Content Creation in the Age of GenAI: To Share or Not to Share?

### v1

@@ C.2 Complete Proof of Theorem 6
Part 2: Computational complexity.
@@
For computing the approximate ESE (Line 5), we leverage Proposition 15, which states that under the assumption $\forall i\in[n],x_{i}\in[0,x_{i}^{\max}]:c^{\prime\prime}_{i}(x_{i})\geq\beta$, the enforced sharing game is strongly monotone. For strongly monotone games, the Multi-Agent Mirror Descent (MAMD) algorithm with full gradient feedback converges at a rate of $O(1/T)$, as established in [12] (see Algorithm 3 for a full specification of the algorithm). Therefore, to find a $\delta$-approximation of an ESE (w.r.t. the $l_{2}$ norm), it suffices to run MAMD for $T=O(1/\delta)$ iterations.
@@
For the $\eta$-FSE check (Line 6), we need to verify if any player can profitably deviate from the profile $(\mathbf{x},\mathbf{1})$ by more than $\eta$. By Proposition 2, for any fixed quality choice $x_{i}$, the optimal sharing level is either $s_{i}=0$ or $s_{i}=1$. Therefore, for each player $i$, we need to check two types of deviations: 1) Deviations where $s_{i}=1$ (changing only quality); and 2) Deviations where $s_{i}=0$ (changing quality and not sharing).
From the proof of Proposition 20, we know that if $\mathbf{x}$ is $\delta$ close to the ESE, then the maximal deviation gain by only varying the quality is at most $\delta M_{1}$. To see why, think of a similar maximal gain from a deviations function that only allows for deviations in $x_{i}$ while keeping the sharing level $s_{i}=1$. As we chose $\delta$ such that $M_{1}\delta\leq\frac{1}{4}\varepsilon$, we can ignore this deviation type as it will not be larger than $\eta$. On the other hand, the second deviation type needs to be checked (recall that not every ESE is an FSE; stability w.r.t. sharing is not guaranteed). We check for such deviations in Algorithm 2.
@@
1:$\mathbf{x}$, $\rho$, $\delta$
2:for $i=1\dots n$ do
3:     Define $f_{0}(x^{\prime}_{i})=\frac{\mu(x^{\prime}_{i}+\|\mathbf{x}_{-i}\|_{1})^{\gamma}x^{\prime}_{i}}{x^{\prime}_{i}+(1+\alpha)\|\mathbf{x}_{-i}\|_{1}}-c_{i}(x^{\prime}_{i})$
4:     Compute $L_{0}$ using Lemma 23
5:     Find $\max_{x^{\prime}_{i}\in[0,x_{i}^{\max}]}f_{0}(x^{\prime}_{i})$ using grid search with step size $\frac{\eta}{L_{0}}$
6:     Let $\phi_{0}=\max_{x^{\prime}_{i}}f_{0}(x^{\prime}_{i})-U_{P}(\mathbf{x},\mathbf{1};f_{\rho})$
7:     if $\phi_{i}>2\eta$ then
8:         return False
9:     else
@@
For each player, we perform a grid search over the interval $[0,x_{i}^{\max}]$ to find the best deviation when switching to withholding all data. By Lemma 23, the utility function for $s_{i}=0$ while the rest of the creators fully share is Lipschitz continuous with constant $L_{0}$, given by:
$L_{0}$ $=\mu\Bigg{[}\frac{\gamma(0.5Y_{LB})^{\gamma-1}(x_{i}^{\max})^{2}}{[(1+\alpha)(0.5Y_{LB})]^{2}}+\frac{\gamma(0.5Y_{LB})^{\gamma-1}x_{i}^{\max}(1+\alpha)(1.5Y_{UB})}{[(1+\alpha)(0.5Y_{LB})]^{2}}$
$+\frac{(x_{i}^{\max}+1.5Y_{UB})^{\gamma}x_{i}^{\max}}{[(1+\alpha)(0.5Y_{LB})]^{2}}+\frac{(x_{i}^{\max}+1.5Y_{UB})^{\gamma}(1+\alpha)(1.5Y_{UB})}{[(1+\alpha)(0.5Y_{LB})]^{2}}$
$+\frac{(1.5Y_{UB})^{\gamma}x_{i}^{\max}}{[(1+\alpha)(0.5Y_{LB})]^{2}}\Bigg{]}+c_{i}^{\prime}(x_{i}^{\max}).$
@@
  • No solution that has larger deviation than $3\eta$ will be accepted, as even with the discretization error reducing the score, the found deviation will be still larger than $2\eta$.
  • Solutions that are between $\eta$ and $3\eta$ might be accepted. This does not bother us, as one can use the analysis we did before to show that solutions that are $3\eta$-FSE will induce a solution that is at most $4\eta$-FSE, which is $\varepsilon$-FSE, and thus we are okay with accepting them, although they are not exactly stand the $\eta$-FSE test.
@@
Wrapping up the complexity analysis; the time complexity of the $\eta$-FSE check for all players is $O\left(\frac{n\cdot L_{0}\cdot\max_{i}x_{i}^{\max}}{\eta}\right)$. Combined with the complexity of finding the approximate ESE, the total time complexity per iteration of the main algorithm is:
$O\left(\frac{n}{\delta}+\frac{n\cdot L_{0}\cdot\max_{i}x_{i}^{\max}}{\delta}\right)=O\left(\frac{n\cdot(1+L_{0}\cdot\max_{i}x_{i}^{\max})}{\eta}\right).$
@@
$O\left(\frac{n\cdot(1+L\cdot\max_{i}x_{i}^{\max})}{\delta^{2}}\right).$

### v2

@@ C.5 Proofs of Section 4’s Main Results
Part 2: Computational complexity
@@
For computing the approximate ESE (Line 6), we leverage Proposition 27, which states that under the assumption $\forall i\in[n],x_{i}\in[0,x_{i}^{\max}]:c^{\prime\prime}_{i}(x_{i})\geq\beta$, the enforced sharing game is strongly monotone. For strongly monotone games, the Multi-Agent Mirror Descent (MAMD) algorithm with full gradient feedback converges at a rate of $O(1/T)$, as established in [12] (see Algorithm˜2 for a full specification of the algorithm). Therefore, to find a $\delta$-approximation of an ESE (w.r.t. the $l_{2}$ norm), it suffices to run MAMD for $T=O(1/\delta)$ iterations.
@@
For the $\eta$-FSE check (Line 7), we need to verify if any player can profitably deviate from the profile $(\mathbf{x},\mathbf{1})$ by more than $\eta$. By Proposition 2, for any fixed quality choice $x_{i}$, the optimal sharing level is either $s_{i}=0$ or $s_{i}=1$. Therefore, for each player $i$, we need to check two types of deviations: 1) Deviations where $s_{i}=1$ (changing only quality); and 2) Deviations where $s_{i}=0$ (changing quality and not sharing).
From the proof of Lemma˜24, we know that if $\mathbf{x}$ is $\delta$ close to the ESE, then the maximal deviation gain by only varying the quality is at most $\delta B(1,(5+\alpha)\cdot 2^{6-\gamma},0.5,3,3-\gamma)$. To see why, think of a similar maximal gain from a deviations function that only allows for deviations in $x_{i}$ while keeping the sharing level $s_{i}=1$. As we chose $\delta$ such that $B(1,(5+\alpha)\cdot 2^{6-\gamma},0.5,3,3-\gamma)\delta\leq\frac{1}{4}\varepsilon$, we can ignore this deviation type as it will not be larger than $\eta$. On the other hand, the second deviation type needs to be checked (recall that not every ESE is an FSE; stability w.r.t. sharing is not guaranteed). We check for such deviations in Algorithm˜4.
@@
[…]
2:Boolean indicating FSE stability
3:for $i=1$ to $n$ do
4:  Define $f_{0}(x^{\prime}_{i})=\frac{\mu(x^{\prime}_{i}+\|\mathbf{x}_{-i}\|_{1})^{\gamma}x^{\prime}_{i}}{x^{\prime}_{i}+(1+\alpha)\|\mathbf{x}_{-i}\|_{1}}-c_{i}(x^{\prime}_{i})$
5:  Compute $L_{0}$ using Lemma 26
6:  Find $\max_{x^{\prime}_{i}\in[0,x_{i}^{\max}]}f_{0}(x^{\prime}_{i})$ using grid search with step size $\frac{\eta}{L_{0}}$
7:  Let $\phi_{i}\leftarrow\max_{x^{\prime}_{i}}f_{0}(x^{\prime}_{i})-U_{i}(\mathbf{x},\mathbf{1};f_{\rho})$
8:  if $\phi_{i}>2\eta$ then
9:   return False
@@
For each player, we perform a grid search over the interval $0,x_{i}^{\max}]$ to find the best deviation when switching to withholding all data. By Lemma˜26, the utility function for $s_{i}=0$ while the rest of the creators fully share is Lipschitz continuous with constant $L_{0}=B(1,5\cdot 2^{5-\gamma},0,2,3-\gamma)$. Therefore, we can perform a grid search with step size $\frac{\eta}{L_{0}}$ to find an approximate maximum of the deviation utility function. This ensures that the discretization error is at most $\eta$.
@@
  • No solution that has a larger deviation than $3\eta$ will be accepted, as even with the discretization error reducing the score, the found deviation will still be larger than $2\eta$.
  • Solutions that are between $\eta$ and $3\eta$ might be accepted. This does not bother us, as one can use the analysis we did before to show that solutions that are $3\eta$-FSE will induce a solution that is at most $4\eta$-FSE, which is $\varepsilon$-FSE, and thus we are okay with accepting them, although they do not exactly stand the $\eta$-FSE test.
@@
Wrapping up the complexity analysis; the time complexity of the $\eta$-FSE check for all players is $O\left(\frac{n\cdot L_{0}\cdot x^{\max}}{\eta}\right)$. Combined with the complexity of finding the approximate ESE, the total time complexity per iteration of the main algorithm is:
$O\left(\frac{n}{\delta}+\frac{n\cdot L_{0}\cdot x^{\max}}{\eta}\right)=O\left(\frac{n\cdot(1+L_{0}\cdot x^{\max})}{\delta}\right).$
@@
$O\left(\frac{n\cdot(1+L_{0}\cdot x^{\max})}{\delta^{2}}\right).$